OpenStudy (chrisga):

Determine the solution to f(x) = g(x) using the following system of equations: f(x) = 3x − 23 g(x) = −4.5x + 7 x = 2 x = 4 x = −2 x = −4

OpenStudy (chrisga):

@LynFran

OpenStudy (lynfran):

3x-23=-4.5x+7 3x+4.5x=7+23 7.5x=30 x=30/7.5 x=4

OpenStudy (chrisga):

oh ok Thank you alot i see how to do it

OpenStudy (lynfran):

welcome